### v2

@@ 4 Manifolds of absolutely continuous functions
The goal is make the family $\{(\mathcal{U}_{\eta},\Psi_{\eta}^{-1}):\eta\in AC_{L^{p}}([a,b],N)\}$ an atlas for $AC_{L^{p}}([a,b],N)$ for a smooth manifold structure on $AC_{Lp}([a,b],N)$. We need to show that the charts are compatibles, i.e., the smoothness of the map
$\Lambda_{\xi,\eta}:=\Psi_{\xi}^{-1}\circ\Psi_{\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\xi),\quad\sigma\mapsto\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma).$ (2)
Let $R=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$, let $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...n\}\}$ and $\{(U_{\phi_{i}},\phi_{i}):i\in\{1,...n\}\}$ be charts that verify the definition of absolutely continuous for $\eta$ and $\xi$ respectively. For $\sigma:[a,b]\to TN$, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$. We will study the smoothness of the composition
$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\to\text{Im}(\Phi_{\xi,R})\subset\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\phi_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})\right)_{i=1}^{n}$
that by Lemma 4.4, is equivalent to the smoothness of $\Lambda_{\xi,\eta}$. For each $i\in\{1,...,n\}$ we have
$d\tilde{\varphi}_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi\right),\Sigma\circ T\varphi_{i}^{-1}\circ T\varphi_{i}\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi\right),\Sigma\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta,d\varphi_{i}\circ\sigma_{i}\right)).$
Lets $H_{i}:\phi_{i}(U_{\phi_{i}})\times\varphi_{i}(U_{i})\times E\to E$ defined by
$H_{i}(x,y,z)=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(x\right),\Sigma\circ T\varphi_{i}^{-1}(y,z)).$ (3)
Since the function $H_{i}$ is smooth (because is composition of smooth functions) the map
$AC([t_{i-1},t_{i}],H_{i}):AC_{L^{p}}([t_{i-1},t_{i}],E\times E\times E)\to AC_{L^{p}}([t_{i-1},t_{i}],E)\quad\alpha\mapsto H_{i}\circ\alpha$
is smooth by Lemma 2.16. Doing the identification of products of $AC_{L^{p}}$ spaces (Remark 2.17), if we fix the functions $\phi_{i}\circ\xi\in AC_{L^{p}}([t_{i-1},t_{i}],E)$ and $\varphi_{i}\circ\eta\in AC_{L^{p}}([t_{i-1},t_{i}],E)$, we see that the map
$\Theta_{i}:AC_{L^{p}}([t_{i-1},t_{i}],TU_{i})\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\tau\mapsto H_{i}\circ(\tilde{\varphi}_{i}\circ\xi,\varphi_{i}\circ\eta,d\varphi_{i}\circ\tau)$
@@
Proceding in the same way, using the fact that composition of $\mathbb{K}$-analytics maps is $\mathbb{K}$-analytic and using the analytic version of Lemma 4.4 (see [9]), we can obtain the analogous case.
@@
Let $N$ be smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $1\leq q\leq p\leq\infty$. Then
@@
Let $\eta\in AC_{L^{p}}([a,b],N)$. Let $\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ such that verify the definition of absolutely continuous for $\eta$. Then, by Proposition 1.2, the smoothness of each inclusion in $\eta$ is equivalent to the smoothness of each inclusion map
$\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)\to\prod_{i=1}^{n}AC_{L^{q}}([t_{i-1},t_{i}],N),\quad(\eta)_{i=1}^{n}\mapsto(\eta)_{i=1}^{n}.$
Since each inclusion $AC_{L^{p}}([t_{i-1},t_{i}],N)\to AC_{L^{q}}([t_{i-1},t_{i}],N)$ is continuous linear [24, Remark 3.2], the inclusions are smooth. ∎
@@
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces $E$ which admits a local addition, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. If $f:M\to N$ is a $C^{k+2}$-map, then the map
@@
Let $(\Omega_{M},\Sigma_{M})$ and $(\Omega_{N},\Sigma_{N})$ be local additions on $M$ and $N$ respectively. The map make sense by Lemma 2.23. Let $\eta\in AC_{L^{p}}([a,b],N)$, $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{f\circ\eta},\Psi_{f\circ\eta}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $f\circ\eta\in AC_{L^{p}}([a,b],N)$ respectively, if $\theta_{N}=(\pi_{TN},\Sigma_{N})$ then we define
$F(\sigma):=\Psi_{f\circ\eta}^{-1}\circ AC_{L^{p}}([a,b],f)\circ\Psi_{\eta}=\theta_{N}^{-1}\circ\Big{(}(f\circ\eta),(f\circ\Sigma_{M})\circ\sigma\Big{)}$
@@
Proceeding as the proof of Theorem 3.1, choosing the corresponding partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ and the families of charts $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},{\phi}_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, we denote $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$. We will study the continuity of the map
@@
where $\Phi_{f\circ\eta}$ is the topological embedding as Proposition 3.2. For each $i\in\{1,...,n\}$ and $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$ we have
$d\phi_{i}\circ F(\sigma)$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}f\circ\eta,f\circ\Sigma_{M}\circ\sigma\Big{)}$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}\circ\phi_{i}\circ f\circ\eta,f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta,d\varphi_{i}\circ\sigma_{i}\right)).$
We define the map $H_{i}:\phi_{i}(U_{\phi_{i}})\times\varphi_{i}(U_{\varphi_{i}})\times E_{M}\to E_{N}$ given by
$H_{i}(x,y,z)=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}(x),f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}(y,z)\Big{)}$
which is a $C^{k+2}$-map. Since $\phi_{i}\circ f\circ\eta\in AC_{Lp}([t_{i-1},t_{i}],E_{N})$ and $\varphi_{i}\circ\eta\in AC_{Lp}([t_{i-1},t_{i}],E_{M})$, then the map
$\Theta_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N})\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ f\circ\eta,\varphi_{i}\circ\eta,\tau)$
is $C^{k}$. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{M})$, $d\sigma\mapsto\varphi_{i}\circ\sigma_{[t_{i-1},t_{i}]}$ are continuous linear by definition, rewriting we have
@@
If $f:M\to N$ is a $\mathbb{K}$-analytic map between $\mathbb{K}$-analytic manifolds modeled on sequentially complete locally convex spaces which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$. Then the map
@@
The manifold structures for $AC_{L^{p}}([a,b],N)$ given by different local additions are diffeomorphic. Indeed, since the identity map $\text{id}_{N}:N\to N$ is smooth regardless of the chosen local addition, the map
$AC([a,b],\text{id}_{M}):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],N),\quad\eta\to\text{id}_{M}\circ\eta$
@@
The inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth. Indeed, let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{J(\eta)}^{C},\Psi_{J(\eta)}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $J(\eta)\in C([a,b],N)$ respectively, then
$\Psi_{J(\eta)}^{-1}\circ J\circ\Psi_{\eta}^{-1}(\sigma):\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap J^{-1}(\mathcal{U}_{j(\eta)})\right)\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$
@@
coincides with the manifold structure on $AC_{L^{p}}([a,b],U)$ constructed with the local addition
$\Sigma|_{\Omega\cap TU}:\Omega\cap TU\to U.$
@@
Let $N_{1}$ and $N_{2}$ be smooth manifolds with local addition modeled on a sequentially complete locally convex space which admits a local addition, $p\in[1,\infty]$ and let $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection for $i\in\{1,2\}$. Then the map
@@
By previous remark, if $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{1},\Sigma_{1})$ are the local addition on $N_{1}$ and $N_{2}$ respectively, then we can assume that the local additions on $N_{1}\times N_{2}$ is
@@
is an diffeomorphism of vector spaces, where $\text{q}_{1}$ and $\text{q}_{2}$ are the corresponding projection of the space, we have
@@
It is clear that the map $T$ is well defined and injective. Let denote $\text{Im}(T)$ the image set of the map $T$, then
@@
Let $\gamma:=(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)$. For each $i\in\{1,...,n\}$ let $\Psi_{\gamma_{i}}^{-1}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ be charts of $\gamma_{i}$, then the map
@@
is a chart for $\gamma$. Let $\eta\in AC_{L^{p}}([a,b],N)$ and $\tilde{\eta}=T(\eta)$, then $\Psi_{\tilde{\eta}}\circ T\circ\Psi_{\eta}^{-1}$ is just restriction of the product of restriction of the smooth maps
@@
thus $T$ is a smooth. Let $\tilde{T}:AC_{L^{p}}([a,b],N)\to\text{Im}(T)$, $\eta\mapsto T(\eta)$, then analogously to Proposition 3.9, we can see that $\tilde{T}^{-1}$ is smooth. Hence $T$ it is a diffeomorphism onto its image. Now we will show that the image $\text{Im}(T)$ is a submanifold. Let $\gamma=(\gamma_{i})_{i=1}^{n}\in\text{Im}(T)$ with charts as before, then for each $i\in\{1,...,n-1\}$ and $\xi=\left(\xi_{i}\right)_{j=1}^{n}\in\text{Im}(T)\cap\prod_{j=1}^{n}\mathcal{U}_{j}$ we have
@@
Then $\Psi_{\gamma}^{-1}|_{\mathcal{U}_{\gamma}\cap\text{Im}(T)}:\text{Im}(\Psi)\cap\prod_{i=1}^{n}\mathcal{U}_{i}\to K\cap\prod_{i=1}^{n}\mathcal{V}_{i}$ is a chart of $\text{Im}(T)$, making $\text{Im}(T)$ a smooth submanifold.
@@
Let $N$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex space which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$. Since every continuous linear operator is analytic, the isomorphism
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. If
$g:[c,d]\to[a,b]$ is a map as in Lemma 2.14, then the map
@@
By Proposition 3.10 we know that the map is well defined. Let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta\circ g},\Psi_{\eta\circ g}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\circ g\in AC_{L^{p}}([c,d],N)$ respectively, then we have
@@
for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}(g,N)^{-1}(\mathcal{U}_{\eta\circ g})\right)$.
Let $\alpha=\eta\circ g:[c,d]\to N$ and $\tau=\sigma\circ g:[c,d]\to TN$, then both are absolutely continuous, with $\pi_{TN}\circ\tau=\alpha$, then $\tau\in\Gamma_{AC}(\alpha)$ and
@@
${\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta\circ g)},\quad\sigma\to\sigma\circ g$

### v4

@@ 3 Manifolds of absolutely continuous functions
The goal is to make the family $\{(\mathcal{U}_{\eta},\Psi_{\eta}^{-1}):\eta\in AC_{L^{p}}([a,b],N)\}$ an atlas for $AC_{L^{p}}([a,b],N)$ for a smooth manifold structure on $AC_{Lp}([a,b],N)$. We need to show that the charts are compatible, i.e., the smoothness of the map
$\Lambda_{\xi,\eta}:=\Psi_{\xi}^{-1}\circ\Psi_{\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\xi),\quad\sigma\mapsto\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma)$
for each $\eta,\xi\in AC_{L^{p}}([a,b],N)$ such that the open set
$\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})=\left(\Psi_{\eta}^{C}\right)^{-1}(\mathcal{U}_{\eta}^{C}\cap\mathcal{U}_{\xi}^{C})\cap{\Gamma_{AC}(\eta)}$
is not empty.
Let $R=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$, let $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},\phi_{i}):i\in\{1,...,n\}\}$ be charts that verify the definition of absolute continuity for $\eta$ and $\xi$, respectively. For each $\sigma\in AC_{L^{p}}([a,b],TN)$, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$ and
$\Lambda_{\xi,\eta}(\sigma_{i}):=\Lambda_{\xi,\eta}(\sigma)\big{|}_{[t_{i-1},t_{i}]}.$
We will study the smoothness of the composition
$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$ $\to\text{Im}(\Phi_{\xi,R})\subset\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)$
$\sigma$ $\mapsto\left(d\phi_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})\right)_{i=1}^{n}$
that by Lemma 3.4, is equivalent to the smoothness of $\Lambda_{\xi,\eta}$. For each $i\in\{1,...,n\}$, we denote $\eta_{i}:=\eta|_{[t_{i-1},t_{i}]}$ and $\xi_{i}:=\xi|_{[t_{i-1},t_{i}]}$ and we have
$d{\phi}_{i}\circ\Lambda_{\xi,\eta}|_{[t_{i-1},t_{i}]}(\sigma_{i})$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ(\xi_{i},\Sigma\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi_{i}\right),\Sigma\circ T\varphi_{i}^{-1}\circ T\varphi_{i}\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi_{i}\right),\Sigma\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i}\right)).$
Because all of the functions involved are continuous and have an open domain, also the composition
$H_{i}(x,y,z):=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(x\right),\Sigma\circ T\varphi_{i}^{-1}(y,z))$ (3.1)
has an open domain $\mathcal{O}_{i}$. Hence the map $H_{i}:\mathcal{O}_{i}\to E$ is smooth. By Lemma 2.17, the map
$AC_{L^{p}}([t_{i-1},t_{i}],H_{i}):AC_{L^{p}}([t_{i-1},t_{i}],\mathcal{O}_{i})\to AC_{L^{p}}([t_{i-1},t_{i}],E)\quad\alpha\mapsto H_{i}\circ\alpha$
is smooth. Doing the identification of products of $AC_{L^{p}}$ spaces (Remark 2.18), if we fix the functions $\phi_{i}\circ\xi_{i}$ and $\varphi_{i}\circ\eta_{i}$, we have the continuous linear map
$AC_{L^{p}}([t_{i-1},t_{i}],E)\to AC_{L^{p}}([t_{i-1},t_{i}],E\times E\times E),\quad\tau\mapsto(\phi_{i}\circ\xi_{i},\varphi_{i}\circ\eta_{i},\tau).$
We write $W_{i}$ for the preimage of $AC_{L^{p}}([t_{i-1},t_{i}],\mathcal{O}_{i})$ under this map. Then the map
$\Theta_{i}:W_{i}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ\xi,\varphi_{i}\circ\eta,d\varphi_{i}\circ\tau)$
@@
Proceding in the same way, using the fact that compositions of $\mathbb{K}$-analytic maps are $\mathbb{K}$-analytic and using the analytic version of Lemma 3.4 (see [13]), we obtain the analogous case.
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space which admits a local addition and $1\leq q\leq p\leq\infty$. Then
@@
Let $\eta\in AC_{L^{p}}([a,b],N)$. Let $\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ that verify the definition of absolute continuity for $\eta$.
By [33, Remark 3.2], we know that each inclusion map $AC_{L^{p}}([t_{i-1},t_{i}],E)\to AC_{L^{q}}([t_{i-1},t_{i}],E)$ is continuous linear, hence $AC_{L^{p}}([a,b],N)\subseteq AC_{L^{q}}([a,b],N)$. By Proposition 2.27, the inclusion is smooth since each map
$\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)\to\prod_{i=1}^{n}AC_{L^{q}}([t_{i-1},t_{i}],E),\quad(\eta_{i})_{i=1}^{n}\mapsto(\eta_{i})_{i=1}^{n}.$
is continuous linear. ∎
@@
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. If $f:M\to N$ is a $C^{k+2}$-map, then the map
@@
Let $E_{M}$ and $E_{N}$ be the modeleding space of $M$ and $N$, respectively. Let $(\Omega_{M},\Sigma_{M})$ and $(\Omega_{N},\Sigma_{N})$ be local additions on $M$ and $N$ respectively. The map makes sense by Lemma 2.25. Let $\eta\in AC_{L^{p}}([a,b],M)$, $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{f\circ\eta},\Psi_{f\circ\eta}^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],M)$ and $f\circ\eta\in AC_{L^{p}}([a,b],N)$, respectively. We see that the set
$\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)={\Gamma_{AC}(\eta)}\cap\left(\Psi_{\eta}^{C}\right)^{-1}\left(\mathcal{U}_{\eta}^{C}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta}^{C})\right)$
is open in ${\Gamma_{AC}(\eta)}$. If $\theta_{N}=(\pi_{TN},\Sigma_{N})$, then we define
$F(\sigma):=\Psi_{f\circ\eta}^{-1}\circ AC_{L^{p}}([a,b],f)\circ\Psi_{\eta}(\sigma)=\theta_{N}^{-1}\circ\Big{(}(f\circ\eta),(f\circ\Sigma_{M})\circ\sigma\Big{)}$
@@
Proceeding as the proof of Theorem 3.5, choosing the corresponding partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ and the families of charts $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},{\phi}_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$, $\eta_{i}:=\eta|_{[t_{i-1},t_{i}]}$ and
$F(\sigma)|_{[t_{i-1},t_{i}]}:=F(\sigma_{i}).$
We will study the continuity of the map
@@
where $\Phi_{f\circ\eta}$ is the topological embedding as in Proposition 2.27. For each $i\in\{1,...,n\}$ and $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$ we have
$d\phi_{i}\circ F(\sigma)|_{[t_{i-1},t_{i}]}$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}f\circ\eta_{i},f\circ\Sigma_{M}\circ\sigma_{i}\Big{)}$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}\circ\phi_{i}\circ f\circ\eta_{i},f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i}\right)\Big{)}.$
Because all of the functions involved are continuous and have an open domain, also the composition
$H_{i}(x,y,z):=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}(x),f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}(y,z)\Big{)}$ (3.2)
has an open domain $\mathcal{O}_{i}$ in $E_{N}\times E_{M}\times E_{M}$. Hence the map $H_{i}:\mathcal{O}_{i}\to E_{N}$ is smooth. Fixing the absolutely continuous functions $\phi_{i}\circ f\circ\eta_{i}$ and $\varphi_{i}\circ\eta_{i}$, we consider the continuous linear map
$AC_{L^{p}}([t_{i-1},i],E_{M})\to AC_{L^{p}}([t_{i-1},i],E_{N}\times E_{M}\times E_{M}),\quad\tau\mapsto(\phi_{i}\circ f\circ\eta,\varphi_{i}\circ\eta,\tau)$
and we write $W_{i}$ for the preimage of $AC_{L^{p}}([a,b],\mathcal{O}_{i})$ under this map. Then the map
$\Theta_{i}:W_{i}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N})\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ f\circ\eta_{i},\varphi_{i}\circ\eta_{i},\tau)$
is $C^{k}$. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{M})$, $\sigma\mapsto d\varphi_{i}\circ\sigma_{i}$ are continuous linear by definition, rewriting we have
@@
Let $f:M\to N$ be a $\mathbb{K}$-analytic map between $\mathbb{K}$-analytic manifolds modeled on sequentially complete locally convex spaces which admit $\mathbb{K}$-analytic local additions and $p\in[1,\infty]$. Then the map
@@
The manifold structures for $AC_{L^{p}}([a,b],N)$ given by different local additions coincide. Indeed, since the identity map $\text{id}_{N}:N\to N$ is smooth regardless of the chosen local addition, the map
$AC([a,b],\text{id}_{M}):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],N),\quad\eta\to\text{id}_{N}\circ\eta$
@@
The inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth. Indeed, let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta}^{C},\left(\Psi_{\eta}^{C}\right)^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\in C([a,b],N)$ respectively, then
$\Psi_{\eta}^{-1}\circ J\circ\Psi_{\eta}^{C}(\sigma):\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap J^{-1}(\mathcal{U}_{\eta}^{C})\right)\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$
@@
coincides with the manifold structure on $AC_{L^{p}}([a,b],U)$.
@@
Let $N_{1}$ and $N_{2}$ be smooth manifolds with local addition modeled on a sequentially complete locally convex spaces which admit a local addition, $p\in[1,\infty]$ and let $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection for $i\in\{1,2\}$. Then the map
@@
By the previous remark, if $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{1},\Sigma_{1})$ are the local addition on $N_{1}$ and $N_{2}$ respectively, then we can assume that the local addition on $N_{1}\times N_{2}$ is
@@
is an isomorphism of topological vector spaces, where $\text{q}_{1}$ and $\text{q}_{2}$ are the corresponding projection of the space, we have
@@
It is clear that the map $T$ is well defined and injective. Let $\text{Im}(T)$ be the image of the map $T$. Then
@@
Let $\gamma:=(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)$. For each $i\in\{1,...,n\}$ let $\Psi_{\gamma_{i}}^{-1}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ be charts around $\gamma_{i}$. Then the map
@@
is a chart around $\gamma$. Let $\eta\in AC_{L^{p}}([a,b],N)$ and $\tilde{\eta}=T(\eta)$, then $\Psi_{\tilde{\eta}}^{-1}\circ T\circ\Psi_{\eta}$ is just restriction of the product of the restrictions of the smooth maps
@@
thus $T$ is a smooth. Now we will show that the image $\text{Im}(T)$ is a submanifold.
Let $\gamma=(\gamma_{i})_{i=1}^{n}\in\text{Im}(T)$ with charts as before, then for each $i\in\{1,...,n-1\}$ and
$\xi=\left(\xi_{i}\right)_{j=1}^{n}\in\text{Im}(T)\cap\prod_{j=1}^{n}\mathcal{U}_{j}$ we have
@@
Then $\Psi_{\gamma}^{-1}|_{\mathcal{U}_{\gamma}\cap\text{Im}(T)}:\text{Im}(\Psi)\cap\prod_{i=1}^{n}\mathcal{U}_{i}\to K\cap\prod_{i=1}^{n}\mathcal{V}_{i}$ is a chart of $\text{Im}(T)$, making $\text{Im}(T)$ a smooth submanifold and the map $\tilde{T}:AC_{L^{p}}([a,b],N)\to\text{Im}(T)$, $\eta\mapsto T(\eta)$ a diffeomorphism. ∎
@@
Let $N$ be a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex space which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$. Since every continuous linear operator is analytic, the isomorphism
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. If $g:[c,d]\to[a,b]$ is a map as in Lemma 2.15, then the map
@@
By Proposition 2.35 we know that the map is well defined. Let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta\circ g},\Psi_{\eta\circ g}^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\circ g\in AC_{L^{p}}([c,d],N)$ respectively, then we have
@@
for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}(g,N)^{-1}(\mathcal{U}_{\eta\circ g})\right)$. This set coincides with
$\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap C(g,N)^{-1}(\mathcal{U}_{\eta\circ g}^{C})\right)$
which is open given by the continuity of the map $C(g,N)$. Let $\alpha=\eta\circ g:[c,d]\to N$ and $\tau=\sigma\circ g:[c,d]\to TN$. Then both are absolutely continuous, with $\pi_{TN}\circ\tau=\alpha$, now $\tau\in\Gamma_{AC}(\alpha)$ and
@@
${\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta\circ g)},\quad\sigma\mapsto\sigma\circ g$